OpenStudy (anonymous):

1073 in scientific notation

OpenStudy (cwrw238):

scientific notation:- N.NN.. * 10^n where N.NN.. is a number between 1 and 9.99...

OpenStudy (cwrw238):

eg 102 = 1.02 * 10^2 49 = 4.9 * 10^2

OpenStudy (cwrw238):

* correction 49 = 4.9 * 10^1

OpenStudy (cwrw238):

do you see how to do them?